The following table shows the prices of a sample of Treasury strips. Each strip makes a single payment at maturity. Years toURGENT!!
Answer #1

n-year rate=(Face Value/Price of n year zero coupon bond)^(1/n)-1

1.
=(1/98.152%)^(1/1)-1
=1.882794%

2.
=(1/94.651%)^(1/2)-1
=2.786812%

3.
=(1/90.844%)^(1/3)-1
=3.252660%

4.
=(1/86.780%)^(1/4)-1
=3.608429%

5.
Upward sloping

6.
Yes this is the usual shape